1
A 10.5-kg block of ice has a temperature of -22.1 °c. The pressure is one atmosphere. The block absorbs 6.99 x of heat. What is the final temperature in degrees Celsius of the liquid 10 the tolerance is +/-2% Click if you would like to show work for this question: open Show works
2
media%2Faa9%2Faa959092-3724-4b63-9b55-25
3
media%2F628%2F6287d79c-5060-425a-8ce3-1c
Answer #1

(1) The final temperature in degree celsius of a liquid water will be given as :

we know that, \DeltaQ = mice Cice (Tf - T0)

(6.99 x 106 J) = (10.5 kg) (2093 J/kg 0C) [Tf + (22.1 0C)]

(6.99 x 106 J) = (21976.5 J/0C) [Tf + (22.1 0C)]

Tf + (22.1 0C) = [(6.99 x 106 J) / (21976.5 J/0C)]

Tf = [(318.06 0C) - (22.1 0C)]

Tf = 295.96 0C

Tf\approx 296 0C

(2) The mass of water that vaporizes which will be given as :

(Heat lost by mercury) = (Heat gained by water) + (Heat needed for phase change)

\DeltaQlost = \DeltaQgained + \DeltaQvaporization

mHg CHg\DeltaT = mw Cw\DeltaT + m \DeltaH

(3.96 kg) (140 J/kg 0C) [(204 - 100) 0C] = (0.212 kg) (4186 J/kg 0C) [(100 - 93.1) 0C] + m (333 x 103 kg)

(57657.6 J) - (6123.2 J)] = m (333 x 103 J/kg)

m = (51534.4 J) / (333 x 103 J/kg)

m = 0.154 kg